# April 3, 2025

### v1

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
For a.a. $t\in(0,T)$, and for all $\kappa>0$, any solution $\bm{u}_{\kappa}$ for Problem $\mathcal{P}_{\kappa}(\Omega)$ satisfies the following equation:
(18) $\ddot{\bm{u}}_{\kappa}(t)+(\tilde{\mathcal{A}}\bm{u}_{\kappa})(t)+(\tilde{\mathcal{B}}\dot{\bm{u}}_{\kappa})(t)+\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}(\tilde{\mathcal{N}}\bm{u}_{\kappa})(t)=\bm{f}(t),\quad\textup{ in }\bm{V}^{\ast}(\Omega).$
Since $\bm{V}^{\ast}(\Omega)\hookrightarrow\bm{H}^{-1}(\Omega)$, we can test (18) along any element $\bm{v}\in\bm{H}^{1}_{0}(\Omega)$. Thanks to (1), we obtain that the penalty term vanishes if, for a.a. $t\in(0,T)$, the linear operator $\mathcal{N}\bm{u}_{\kappa}(t)$ is tested along an element $\bm{v}\in\bm{H}^{1}_{0}(\Omega)$. Therefore, the function $\bm{u}_{\kappa}$ solves the following equation
in the sense of distributions in $(0,T)$. Note that (19) is a linear equation; this implies that:
$\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}\textup{ is bounded in }L^{2}(0,T;\bm{H}^{-1}(\Omega))\textup{ independently of }\kappa.$
An application of Theorem 10.1.20 of [34] gives that $\dot{\bm{u}}_{\kappa}\rightharpoonup\dot{\bm{u}}$ in $\mathcal{C}^{0}([0,T];\bm{H}^{-1}(\Omega))$ as $\kappa\to 0$ up to passing to a suitable subsequence, thus verifying the initial condition for the velocity and establishing $(i)$.
To establish that the acceleration $\ddot{\bm{u}}$ enjoys the regularity of a vector-valued measure in the sense of Dinculeanu (cf., e.g., [3, 13]), we first observe that if we test the equation (18) at an arbitrary $\bm{v}\in\bm{V}(\Omega)$ such that $\|\bm{v}\|_{\bm{V}(\Omega)}=1$ and we integrate between 00 and $T$, we obtain that:
(20) $\dfrac{1}{\kappa}\left|\left\langle\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t,\bm{v}\right\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|$
$\dfrac{1}{\kappa}\left|\int_{0}^{T}\chi_{(0,T_{0})}(t)\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right|$
$\leq\int_{0}^{T}\|\mathcal{A}\bm{u}_{\kappa}(t)\|_{\bm{V}^{\ast}(\Omega)}+\|\mathcal{B}\dot{\bm{u}}_{\kappa}(t)\|_{\bm{V}^{\ast}(\Omega)}\,\mathrm{d}t+\|\dot{\bm{u}}_{\kappa}(T)-\bm{u}_{1}\|_{\bm{V}^{\ast}(\Omega)}.$
We observe that the right-hand side of (20) is bounded independently of $\kappa$ in light of Lemmas 4.24.4, thus showing that:
$\left\{\left\|\dfrac{1}{\kappa}\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t\right\|_{\bm{V}^{\ast}(\Omega)}\right\}_{\kappa>0}\textup{ is bounded independently of }\kappa.$
By definition of supremum, for a.a. $t\in(0,T)$, we can find a sequence $\{\bm{v}_{t}^{(k)}\}_{k=1}^{\infty}\subset\bm{V}(\Omega)$ such that $\|\bm{v}_{t}^{(k)}\|_{\bm{V}(\Omega)}=1$, for which,
$\lim_{k\to\infty}\left|\langle\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}_{t}^{(k)}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|=\sup_{\begin{subarray}{c}\bm{v}\in\bm{V}(\Omega)\\ \|\bm{v}\|_{\bm{V}(\Omega)=1}\end{subarray}}\left|\langle\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|,$
and the supremum on the right-hand side of the latter is realised by the vector fields $\bm{v}\in\bm{V}(\Omega)$ such that $\|\bm{v}\|_{\bm{V}(\Omega)}=1$ such that $(v_{i}\bm{e}^{i}\cdot\bm{q})$ has negative sign a.e. in $\Omega$. In correspondence of such elements, we can drop the absolute value. Keeping the latter statement in mind, and applying the Fatou lemma, we obtain that:
$\int_{0}^{T}\left\|\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\right\|_{\bm{V}^{\ast}(\Omega)}\,\mathrm{d}t\leq\liminf_{k\to\infty}\int_{0}^{T}\chi_{(0,T_{0})}(t)\left|\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}_{t}^{(k)}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|\,\mathrm{d}t$
$\leq\sup_{\begin{subarray}{c}\bm{v}\in\bm{V}(\Omega)\\ \|\bm{v}\|_{\bm{V}(\Omega)=1}\end{subarray}}\int_{0}^{T}\chi_{(0,T_{0})}(t)\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t$
$\leq\sup_{\begin{subarray}{c}\bm{v}\in\bm{V}(\Omega)\\ \|\bm{v}\|_{\bm{V}(\Omega)=1}\end{subarray}}\left|\left\langle\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t,\bm{v}\right\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|=\left\|\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t\right\|_{\bm{V}^{\ast}(\Omega)}.$
In conclusion, we have obtained that
$\int_{0}^{T}\left\|\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\right\|_{\bm{V}^{\ast}(\Omega)}\,\mathrm{d}t=\left\|\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t\right\|_{\bm{V}^{\ast}(\Omega)},$
and that:
(21) $\left\{\dfrac{1}{\kappa}\chi_{(0,T_{0})}\tilde{\mathcal{N}}\bm{u}_{\kappa}\right\}_{\kappa>0}\textup{ is bounded in }L^{1}(0,T;\bm{V}^{\ast}(\Omega))\textup{ independently of }\kappa.$
In light of Lemmas 4.24.4 (in particular, here we need the boundedness of the sequence $\{\dot{\bm{u}}_{\kappa}\}_{\kappa>0}$ in $L^{2}(0,T;\bm{V}(\Omega))$ independent of $\kappa$) and (21), we obtain that:
$\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}\textup{ is bounded in }L^{1}(0,T;\bm{V}^{\ast}(\Omega))\textup{ independently of }\kappa.$
The latter statement puts us in position to apply Theorem 4.3 so that, up to passing to subsequences, we obtain:
$\ddot{\bm{u}}_{\kappa}\overset{\ast}{\rightharpoonup}\bm{\mu},\quad\textup{ in}\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega))\textup{ as }\kappa\to 0.$
Observe that the vector-valued measure $\bm{\mu}\in\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega))$ can be understood as the acceleration of $\bm{u}$. This interpretation arises from the classical definition of the weak derivative, which states that:
$\int_{0}^{T}\dot{\bm{u}}_{\kappa}(t)\varphi^{\prime}(t)\,\mathrm{d}t=-\int_{0}^{T}\ddot{\bm{u}}_{\kappa}(t)\varphi(t)\,\mathrm{d}t,\quad\textup{ for all }\varphi\in\mathcal{D}(0,T).$
For all $\bm{v}\in\bm{V}(\Omega)$ and all $\varphi\in\mathcal{D}(0,T)$, it results
$\int_{\Omega}\int_{0}^{T}\dot{u}_{i,\kappa}(t)(\varphi^{\prime}(t)v_{i})\,\mathrm{d}t\,\mathrm{d}x=-\int_{\Omega}\int_{0}^{T}\ddot{u}_{i,\kappa}(t)(\varphi(t)v_{i})\,\mathrm{d}t\,\mathrm{d}x,$
so that, letting $\kappa\to 0$ (see Comment 3 of Chapter 4 of [5]) gives
$\int_{\Omega}\left(\int_{0}^{T}\dot{u}_{i}(t)\varphi^{\prime}(t)\,\mathrm{d}t\right)v_{i}\,\mathrm{d}x=\int_{0}^{T}\left(\int_{\Omega}\dot{u}_{i}(t)v_{i}\,\mathrm{d}x\right)\varphi^{\prime}(t)\,\mathrm{d}t=-\int_{0}^{T}\int_{\Omega}\,\mathrm{d}\mu_{i}(t)(v_{i}\varphi(t))\,\mathrm{d}t,$
where the first equality holds by Fubini’s theorem (cf., e.g., [29]). To sum up, we have obtained that
$\int_{0}^{T}\int_{\Omega}\dot{u}_{i}(t)v_{i}\,\mathrm{d}x\varphi^{\prime}(t)\,\mathrm{d}t=-\int_{0}^{T}\int_{\Omega}\,\mathrm{d}\mu_{i}(t)v_{i}\varphi(t)\,\mathrm{d}t,$
for all $\varphi\in\mathcal{D}(0,T)$ and all $\bm{v}\in\bm{V}(\Omega)$. We can therefore interpret the vector-valued measure $\bm{\mu}$ as the acceleration $\bm{u}$. This allows us to adopt the notation $\bm{\mu}=\ddot{\bm{u}}$, and the proof of $(ii)$ is thus complete. ∎
We now observe that further compactness can be extracted in light of Lemma 4.4 and Lemma 4.5.
Lemma 4.6.
Consider the sequence $\{\bm{u}_{\kappa}\}_{\kappa>0}$, where each $\bm{u}_{\kappa}$ is a solution of Problem $\mathcal{P}_{\kappa}(\Omega)$. Up to passing to a subsequence, the following convergences hold:
$\dot{\bm{u}}_{\kappa}\to\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Omega))\textup{ as }\kappa\to 0.$
Proof.
Recall that in Lemma 4.5, we have shown that $\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}$ is bounded in $L^{1}(0,T;\bm{V}^{\ast}(\Omega))$ independently of $\kappa$. Thanks to (8), we are allowed to apply the Aubin-Lions-Simon theorem (cf., e.g., Theorem 8.62 in [29]) gives that there exists a subsequence, still denoted $\{\dot{\bm{u}}_{\kappa}\}_{\kappa>0}$ such that
$\dot{\bm{u}}_{\kappa}\to\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Omega))\textup{ as }\kappa\to 0,$
as it had to be proved. ∎
We are now ready to establish the existence of solutions for Problem $\mathcal{P}(\Omega)$.
Theorem 4.4.
Assume that the applied body force $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$ is admissible in the sense of Definition 3.1. Then, Problem $\mathcal{P}(\Omega)$ admits at least one solution.
Proof.
Let $\bm{v}\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$. Let us test equation (18) at $(\bm{v}(t)-\bm{u}_{\kappa}(t))\in\bm{V}(\Omega)$ for all $t\in[0,T]$. In light of Lemma 4.5, it is straightforward to see that:
(22) $\int_{0}^{T}\langle\ddot{\bm{u}}_{\kappa}(t),\bm{v}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\to\langle\langle\ddot{\bm{u}},\bm{v}\rangle\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)},\quad\textup{ as }\kappa\to 0.$
An application of Lemma 4.6 gives:
(23) $\limsup_{\kappa\to 0}\left(-\int_{0}^{T}\langle\ddot{\bm{u}}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)$
$=-\lim_{\kappa\to 0}\langle\dot{\bm{u}}_{\kappa}(T),\bm{u}_{\kappa}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+\limsup_{\kappa\to 0}\|\dot{\bm{u}}_{\kappa}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}$
$=-\langle\dot{\bm{u}}(T),\bm{u}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+\|\dot{\bm{u}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}.$
Thanks to Lemma 4.4, we infer that the convergence $\bm{u}_{\kappa}(T)\rightharpoonup\bm{u}(T)$ in $\bm{V}^{\ast}(\Omega)$ as $\kappa\to 0$ holds. This implies that:
(24) $\limsup_{\kappa\to 0}\left(-\int_{0}^{T}\langle\mathcal{B}\dot{\bm{u}}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)\leq\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x$
$\quad-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(T))e_{i\|j}(\bm{u}_{\kappa}(T))\,\mathrm{d}x.$
In a similar way, we obtain that the convergence $\bm{u}_{\kappa}\rightharpoonup\bm{u}$ in $L^{2}(0,T;\bm{V}(\Omega))$ as $\kappa\to 0$ implies that:
(25) $\limsup_{\kappa\to 0}\left(-\int_{0}^{T}\langle\mathcal{A}\bm{u}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)\leq-\int_{0}^{T}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}(t))e_{i\|j}(\bm{u}(t))\,\mathrm{d}x\,\mathrm{d}t.$
For what concerns the penalty term, we observe that
@@
(26) $-\dfrac{1}{\kappa}\int_{0}^{T}\chi_{(0,T_{0})}(t)\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\left([\bm{I}+v_{i}(t)\bm{e}^{i}]\cdot\bm{q}-[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right)\,\mathrm{d}\Gamma\,\mathrm{d}t\leq 0,$
for all $\kappa>0$.
$2\rho\langle\langle\ddot{\bm{u}},\bm{v}\rangle\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}-2\rho\langle\dot{\bm{u}}(T),\bm{u}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+2\rho\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+2\rho\|\dot{\bm{u}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}$

### v3

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
Let $\kappa>0$ be given. To show that the function $E_{\kappa}$ is of class $L^{1}(T_{0},T)$ we proceed as in Lemma 4.6. Observe that, thanks to Corollary 10.1.26 of [43], it results that $E_{\kappa}$ is differentiable a.e. in $(T_{0},T)$ and that $\dot{E}_{\kappa}\in L^{1}(T_{0},T)$.
Differentiating $E_{\kappa}$ and invoking the variational equations of Problem $\mathcal{P}_{\kappa}(\Omega)$ tested at $\bm{v}=\bm{u}_{\kappa}(t)$ give:
(43) $\dot{E}_{\kappa}(t)=\rho\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Omega}|\dot{\bm{u}}_{\kappa}(t)|^{2}\,\mathrm{d}x\right)+\dfrac{1}{2}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{u}_{\kappa}(t))\,\mathrm{d}x\right)$
@@
$\quad+\dfrac{1}{2\kappa}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Gamma}\left|\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\right|^{2}\,\mathrm{d}\Gamma\right)$
$=-\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}_{\kappa}(t))\,\mathrm{d}x\leq 0,$
showing that $\dot{E}_{\kappa}\in L^{1}(T_{0},T)$. Therefore,in the same spirit as Lemma 4.6, we obtain that $E_{\kappa}\in W^{1,1}(T_{0},T)\subset AC([T_{0},T])$ and that $E_{\kappa}$ decreases in $T_{0},T]$. The continuity of $E_{\kappa}$ in $T_{0},T]$ implies that the mapping
(44) $0,T]\ni t\mapsto\int_{\Omega}|\dot{\bm{u}}_{\kappa}(t)|^{2}\,\mathrm{d}x,$
is continuous and, thanks to (24), is bounded in $T_{0},T]$ independently of $\kappa$.
In correspondence of $\varepsilon>0$, that we will determine later, we denote and define the modified energy $H_{\kappa,\varepsilon}$ by:
$H_{\kappa,\varepsilon}(t):=E_{\kappa}(t)+\varepsilon\rho\int_{\Omega}\dot{\bm{u}}_{\kappa}(t)\cdot\bm{u}_{\kappa}(t)\,\mathrm{d}x,\quad\textup{ for a.a. }t\in(T_{0},T).$
Thanks to the regularity of $\bm{u}_{\kappa}$ in Lemma 4.6 and thanks to the regularity of $E_{\kappa}$ observed beforehand, we can differentiate $H_{\kappa,\varepsilon}$ with respect to the variable $t$, getting
(45) $\dot{H}_{\kappa,\varepsilon}(t)=\dot{E}_{\kappa}(t)+\varepsilon\rho\langle\ddot{\bm{u}}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+\varepsilon\rho\int_{\Omega}|\dot{\bm{u}}_{\kappa}(t)|^{2}\,\mathrm{d}x$
$=-\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}_{\kappa}(t))\,\mathrm{d}x-\dfrac{\varepsilon}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{u}_{\kappa}(t))\,\mathrm{d}x$